OpenStudy (anonymous):

What is the wavelength (in nm) of radiation that has an energy content of 1.0x10^2 kJ/mol? In which region of the electromagnetic spectrum is this radiation found? The answer is 1.2 * 10 ^ 2 nm but I cannot figure it out

OpenStudy (anonymous):

Energy = h nu, where nu is the frequency h = 6.63 x 10^-34 J-s, Planck's constant So nu = E/h = 1 x 10^5 J /h = 0.15 x 10^29 / s nu lambda = c, the speed of light. lambda = wavelength = c / nu =3 x 10^8 / 0.15 x 10^29 = 20 x 10^-21 m. Very, very short wavelength.. Gamma or x-rays?